AnakKrakatau1883:

Harry had $32. He spent all the money buying three notebooks for x dollars each and four packs of index cards for y dollars each. If Harry had bought five notebooks and five packs of index cards, he would have run short by $18. The following system of equations models this scenario: 3x + 4y = 32 5x + 5y = 50 Use the system of equations to solve for x and y. (8, 2) (1, 5) (2, 8) (5, 1)

CheeseStick7793:

3x+4y=32 ( here is spent all money so I add 32 here ) 5x+5y=50 by elemination method 3x+4y=32×5 5x+5y=50×4 ( here I'm going to eliminate y ) 15 x+20y= 160 20x+20y= 200 - - = - ____________ -5x = -40 x = - 40 / -5 x = 8 here - × - = + so iam taking positive 8 now substitute the value of x in any equation 3×8+4y = 32 24 + 4y = 32 4y = 32-24 y = 8 / 4 y = 2 CHECKING SUBSTITUTE BOTH VALUES IN ANY EQUATION 5×8+5×2 = 50 40 + 10 = 50 50 = 50

CheeseStick7793:

x = 8 y = 2